What is the molarity of a solution made by dissolving 5.25 g of potassium
hydrogen phthalate in distilled water and making it up to 250 mL with
distilled water in a volumetric flask?
1
Expert's answer
2019-09-04T03:21:40-0400

First of all find the amount of potassium hydrogen phthalate:


"\nu = m/M"

Find molar mass of potassium hydrogen phthalate (KHC8H4O4):


"M = 39 + 1 + (8*12) + 4 + (4*16) = 204 (g\/mol)"

"\nu = 5.25/204 = 0.0257 (mol)"


Find the molarity of the solution, if there is 0.0257 mol in 250 ml:

"C_m = \nu/V"

"C_m = 0.0257\/0.25 = 0.1028 (mol\/L)"


Answer: the molarity of the potassium hydrogen phthalate solution is 0.1028 mol/L.